A basket has 12 marigolds and 8 carnations. if we had picked two flowers one by one with replacement, find the probability of getting a marigold and a carnation.


1
Expert's answer
2022-04-27T01:06:07-0400
"P(MC)=\dfrac{12}{12+8}(\dfrac{8}{12+8})=0.24"

"P(CM)=\dfrac{8}{12+8}(\dfrac{12}{12+8})=0.24"

"P(M\ and \ C)=P(MC)+P(CM)"

"=0.24+0.24=0.48"

The probability of getting a marigold and a carnation is 0.48.